name a molecule which contains both a single bond and a double bond between carbon and oxygen atom.
1
Expert's answer
2013-02-15T05:59:30-0500
There are several types of molecules possible which satisfy such description:

Carboxylic acids:

R-C=O
I
OH

Esthers:

R-C=O
I
O-R

Carbonates:

Me-O-C=O
I
O
I
Me